OpenStudy (anonymous):

7 − (5 − 8x) = 4x + 3 anyone help with distributive properties i get this far, then blank out.. 35-56x =4x+3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[7-(5-8x)=4x+3\] \[7-5+8x=4x+3\] \[2+8x=4x+3\] \[8x-4x=3-1\] \[4x=2\] \[x=\frac{2}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

step 1, remove parentheses distributing the minus sign. step 2, combine like terms 7-5=2 step 3, subtract 4x from both sides and subtract 2 from both sides step 4 , divide by 2

OpenStudy (anonymous):

oops i made a mistake! \[2+8x=4x+3\] \[8x-4x=3-1\] \[4x=1\] \[x=\frac{1}{4}\]

OpenStudy (anonymous):

for some reason i cannot type a 2 is is 8x-4x=3-2

OpenStudy (anonymous):

4x=1 x=1/4

OpenStudy (anonymous):

i have that as the answer 1/4, i just get hung up on d prop thank you so much...

OpenStudy (anonymous):

welcome! sorry about the error

OpenStudy (anonymous):

i thought i multiplied 7 & 5 then 7 & 8 when distributing the prop